OpenStudy (anonymous):

5(8 - n) = 3n-16 Please show me the steps on how to solve thanks

OpenStudy (ajprincess):

using distributive property distribute the expression given in left hand side. Distributive property is a(b+c)=a*b+a*c

OpenStudy (anonymous):

neeed to know how to do that step by step

OpenStudy (ajprincess):

the expression given in left hand side is 5(8-n). Using distributive property we can distribute it as 5(8-n)=5*8-5*n =40-5n. getting it @darkside3704?

OpenStudy (ajprincess):

no

OpenStudy (anonymous):

yeah, you might want to type out the steps fully so i can get it. other wise i am lost